Find the fourier transform of the function below, considering a, b and c as constants.

50.5 160- 10 -> Other
Answer #1

Fourier trens form of segmeed Hit) is given cry Xcw) - Priete justato Xefir net ë draft at o ifret) = 56+) +şc2-5a4? - -usstswt 0200 Fre) = 1 btl. 5 2 = C - 5a12d 3 dt. 1. . 3 1 to 209 и te dt *507 to jut dt - 5a 12-jut - dt. 31 c -0-5 WI- 4 ell -0.5Iza 5c² r -jwa/ jw I eae (-3 $w) ji e) I a=5 c²r Izjak => I 2 =10c² - 3W a long time sinifwa [one e pres] e-pto 7Part 1 put sa-jw In- saytestant - Integrating by partsy 3 112 atin F-112 S2 as I = 5b ftes st 712 C put sa =) = 5b rejurar ju( 1010. ponose I = 5 b xj [ Coswt) at 2 Sin (with) 1 . -jw r3 JL sa 56 (w cela, 3 - 2 ] let us salvei part 3 monu now ,+ L -12 Usiry Integration by parts, 12250 ,2- S Lost -te -12 tedt To se forman la 5350 spezielle font] . Again using integrat. a) Iz = 5a linaw Le justo I Cewth just 3jw| - w2 jwt -jwt 7 + 7 / 2 e dx2 + le 2 te as Izaba 3iw 1 4 zi jwx 2 jwth jwtra] -1. Let us Summa rise. ful frutta IN I Fiw) - Ipt. Iz t Iz. Pluie Die beste werkg)-2 since] Izz todo sin(wa) Is C e suley-w-so